# Towards General Industrial Intelligence: A Survey on Industrial IoT-Enhanced Continual Large Models

### v1

@@ Abstract
Currently, most applications in the Industrial Internet of Things (IIoT) still rely on CNN-based neural networks. Although Transformer-based large models (LMs), including language, vision, and multimodal models, have demonstrated impressive capabilities in AI-generated content (AIGC), their application in industrial domains, such as detection, planning, and control, remains relatively limited. Deploying pre-trained LMs in industrial environments often encounters the challenge of stability and plasticity due to the complexity of tasks, the diversity of data, and the dynamic nature of user demands. To address these challenges, the pre-training and fine-tuning strategy, coupled with continual learning, has proven to be an effective solution, enabling models to adapt to dynamic demands while continuously optimizing their inference and decision-making capabilities. This paper surveys the integration of LMs into IIoT-enhanced General Industrial Intelligence (GII), focusing on two key areas: LMs for GII and LMs on GII. The former focuses on leveraging LMs to provide optimized solutions for industrial application challenges, while the latter investigates continuous optimization of LMs learning and inference capabilities in collaborative scenarios involving industrial devices, edge computing, and cloud computing. This paper provides insights into the future development of GII, aiming to establish a comprehensive theoretical framework and research direction for GII, thereby advancing GII towards a more general and adaptive future.
@@
General Industrial Intelligence, Large Language Models, Continual Learning, Industrial Internet of Things, Edge Computing.
@@ I Introduction
Industrial 5.0 integrates cyber-physical-social elements into manufacturing, emphasizing the interaction between digital and physical systems as well as human-machine collaboration, effectively connecting devices, objects, and people through the Internet [1]. With the rapid development of the Internet of Things (IIoT) [2, 3, 4], communication technologies [5, 6], AI-generated content (AIGC) [7], robotics and digtal twins [8, 9, 10], modern industrial systems are becoming increasingly complex. These systems not only generate high-frequency unimodal data but also encompass multimodal data types such as text, images, videos, code, and audio. Industrial big data can be leveraged to create digital manufacturing workflows and industrial processes, significantly driving improvements in productivity, efficiency, and effectiveness within Industry 5.0 and cyber-physical-social systems.
@@
As illustrated in Fig. 1, datasets and models constitute the foundational elements of the GII ecosystem, driving the rapid development of higher-level algorithms and applications. These applications, which include intelligent control systems, predictive maintenance [11], fault diagnosis [12, 13], and anomaly detection [14], rely heavily on the extraction and analysis of IIoT data. The success of GII is particularly dependent on its ability to efficiently learn and extract valuable features from these IIoT datasets.
Large models (LMs) based on the Transformer [15], such as Large Language Models (LLMs) [16, 17, 18], vision models [19, 20], time series models [21], and multimodal models [22, 23], have gained widespread attention due to their unique advantages. With billions to trillions of parameters, these models accumulate extensive knowledge through pre-training on large datasets, significantly advancing the automation and diversification of data processing while reducing reliance on human expertise.
In the industrial sector, the precision and scalability of LMs make them highly effective for enhancing the accuracy of industrial processes. However, deploying pre-trained LMs in industrial settings requires careful adaptation to specific task architectures, dynamic data distributions, and user preferences. Balancing stability and adaptability in these environments remains a significant challenge, despite the advantages of LMs in multitask generalization, few-shot learning, and inference.
Inspired by the success of LMs in natural language processing (NLP), the industrial community has increasingly explored their potential for GII. One approach involves building industry-specific foundational models from scratch [24], though the limited scale of domain-specific data often hinders the development of emergent capabilities. Alternatively, pre-training on large datasets followed by task-specific fine-tuning has shown immense potential for building robust industrial models, significantly improving performance across diverse tasks. This method effectively addresses the challenges posed by data scarcity in specific domains, while accelerating the development of advanced capabilities in industrial applications.
Adapting LMs for industrial tasks represents a significant research direction [25]. The advantages of these models in cross-task generalization, zero-shot/few-shot learning, and inferencing capabilities offer new avenues for addressing challenges in knowledge transfer, data scarcity, and explainability in GII. Continual Large Models (CLMs) play a crucial role in maintaining and evolving the capabilities of these industrial models over time. Pre-trained on extensive datasets and powered by Transformer architectures, CLMs are designed to continuously learn and adapt, ensuring that industrial large models remain relevant and effective in meeting the evolving demands of GII.
I-A Goals of This Survey
This paper aims to establish a comprehensive perspective and conduct an in-depth analysis of IIoT-Enhanced GII. It proposes a bifurcation of GII into two main categories:
$\bullet$ Large Models for General Industrial Intelligence (LMs for GII): This direction focuses on utilizing the advanced data processing and analysis capabilities of LMs to solve optimization problems inherent in industrial applications. Specifically, LMs enhance the intelligence and operational efficiency of IIoT-driven industrial systems by leveraging their ability to process real-time multimodal IIoT data, perform intricate feature extraction, and ensure precise pattern recognition and result validation, ultimately improving measurement accuracy and system performance across diverse industrial settings.
$\bullet$ Large Models on General Industrial Intelligence (LMs on GII): This perspective examines how industrial applications can extend and refine the capabilities of LMs through continuous model operation in collaborative IIoT device-edge-cloud environments. By adopting continual learning (CL) and online learning strategies, models can adapt to new data and environmental changes without the need for costly retraining. This approach conserves computational resources, minimizes latency, and efficiently handles distribution shifts and performance degradation, ensuring robust model performance in dynamic industrial scenarios.
The introduction of these categories is further clarified with an illustration (Fig. 2), which helps delineate the structural differences and operational mechanisms between the two approaches.
Figure 2: A high-level overview of LMs for GII and LMs on GII. The vertical dimension primarily focuses on leveraging the general knowledge of LMs to enhance General Industrial Intelligence. The horizontal dimension concentrates on the continual updating of LMs in long sequence industrial tasks.
Large Models (LMs) with General Industrial Intelligence (GII)LMs for GII (§$\S$§III)Preliminary (§$\S$§II)LMs on GII (§$\S$§IV)Outlook and Future Opportunities (§$\S$§V)Research Roadmap (§$\S$§II-A)Industrial LMs (§$\S$§II-B)Key Tasks of GII (§$\S$§II-C) Classification (§$\S$§II-C1), Detection/Localization (§$\S$§II-C2), Planning (§$\S$§II-C3), General Q&A (§$\S$§II-C4) LMs as a Data Enhancer (§$\S$§III-A)LMs as a Tuning Predictor (§$\S$§III-B)LMs as a Non-Tuning Predictor (§$\S$§III-C)Classical Continual Learning (§$\S$§IV-A)Pre-trained Models Based CL (§$\S$§IV-B)Efficient Inference Algorithm (§$\S$§IV-C)Benchmarks (§$\S$§IV-D) Model Architectures (§$\S$§II-B1), LLMs (§$\S$§II-B2), Vision-based Models (§$\S$§II-B3), Time Series-based Models (§$\S$§II-B4), Multimodal Models (§$\S$§II-B5) Data Denoise (§$\S$§III-A1), Data Selection (§$\S$§III-A2), Data Augmentation (§$\S$§III-A3), Data Generation (§$\S$§III-A4), Reprogramming (§$\S$§III-A5) Alignment Fine-tuning (§$\S$§III-B1), Efficient Tuning (§$\S$§III-B2) Regularization-based Methods (§$\S$§IV-A1), Replay-based Methods (§$\S$§IV-A2), Parameter Isolation Methods (§$\S$§IV-A3) Mixture of Experts (§$\S$§IV-B1), Analytic Learning-based Methods (§$\S$§IV-B2), Prompt-based Methods (§$\S$§IV-B3), Representation-based Methods (§$\S$§IV-B4) Test Time Adaptation (§$\S$§IV-C1), Edge Cloud Collaboration (§$\S$§IV-C2), Federated Learning (§$\S$§IV-C3) Data Generation (§$\S$§V-A), Model/Knowledge Reuse (§$\S$§V-B), Beyond Single Modality (§$\S$§V-C), Learning with Restricted Resources (§$\S$§V-D)
Figure 3: Organization of our survey.
I-B Distinguishing Features of This Survey
CL has gained significant traction as a research topic in recent years, with numerous studies exploring its application across various domains such as equipment health management [26], robotics [27], and streaming data [28]. In the context of LMs, CL has been recognized as crucial due to the impracticality of frequent retraining caused by the high costs associated with these models’ massive scale [29].
Despite the extensive literature on CL, our survey uniquely addresses the continual adaptation of LMs for IIoT-enhanced industrial systems—a field not thoroughly covered by existing literature. This survey is the pioneer in providing a comprehensive and systematic review of CL methods tailored for LMs across four distinct IIoT-enabled industrial scenarios.
TABLE I: Overview of General Industrial Intelligence Perspectives in Different Surveys (based on the hierarchical structure shown in Fig. 1).
Layer Perspectives Surveys
Application Predictive Maintenance Hurtado et al. 2023 [26]
Application Industrial Detection Li et al. 2024 [30]
Application Autonomous Driving Katare et al. 2023 [31]
Application AIGC Xu et al. 2024 [7]
Algorithm Continual Learning Ours
Foundation Model Design Liu et al. 2022 [32]
Foundation Federated Learning Nguyen et al. 2021 [33]
Infrastructure End-Edge-Cloud Wang et al. 2024 [34]
Infrastructure Green AI Mao et al. 2023 [35]
Else Safety Perez et al. 2023 [36]
Outlined in Table I, this survey distinguishes itself through several key contributions:
$\bullet$ Novel Taxonomy: We introduce a novel theoretical framework for GII. By categorizing the application of large models into two dimensions—“LMs for GII” and “LMs on GII”—this paper not only explores how to optimize industrial applications using large models but also investigates how these applications can in turn refine the models themselves. This bidirectional interaction perspective significantly enriches the existing literature.
$\bullet$ Cross-Domain Multimodal Integration: Unlike most existing studies that focus solely on specific types of large models, such as language models or vision models, this review encompasses LLMs, vision transformers, multimodal models, and time series models. This cross-modal integration enhances the design, development, and evaluation of complex instrumentation and measurement systems used in generating, acquiring, conditioning, and processing signals. By leveraging the unique advantages across different models, it offers a more comprehensive and in-depth perspective for advancing measurement science and applications, thereby addressing complex industrial challenges more effectively.
$\bullet$ Practical Applications of Continual Learning: This paper emphasizes the real-world application of CL strategies in IIoT-enabled industrial systems, particularly in edge and cloud computing collaborative environments. This focus ensures that the models are not only adaptable to new data and changing conditions but are also resource-efficient. By reducing computational demands and training costs, our approach addresses critical constraints in industrial applications.
@@ II-A Research Roadmap
$\bullet$ CLMs, where models can continuously learn in changing tasks and environments, significantly expanding the boundaries of machine learning applications, marking a step towards higher levels of GII. Representative models include GPT [18] and LLaMA [16].
@@
II-B Industrial Large Models
Industrial LMs are large foundational models pre-trained on diverse industrial data sources, covering various data granularities and modalities, as depicted in Fig. 5.
@@ II-B1 Model Architectures
Figure 5: Industrial LMs are pre-trained on multi-source, multi-granularity and multimodal industrial data and can be adapted for a variety of downstream applications.
@@ II-B5 Multimodal Models
Multimodal models typically integrate various types of inputs, such as text, images, and audio, within a single framework for data processing and analysis. By harnessing the fusion of multiple perceptual information, these models aim to enhance task performance accuracy and efficiency.
Multimodal Large Language Models (MLLMs) combine the linguistic processing capabilities of LLMs with functionalities for processing other modalities, making them particularly adept at understanding and generating content involving complex data types, such as text and images. Notable examples include:
$\bullet$ BLIP [48] is designed to enhance image description and comprehension. Utilizing pre-training techniques such as contrastive learning. BLIP-2 [49] employs a dual-stream Transformer structure that optimizes the processing of visual and textual streams. With an improved cross-modal attention mechanism, BLIP-2 achieves more precise alignment between images and text, enhancing the model’s ability to describe and understand image content.
$\bullet$ MiniGPT-4 [50], a scaled-down version of the GPT, focuses on small-scale but efficient image and text processing tasks, maintaining the generative capabilities and flexibility of the GPT series.
$\bullet$ LLaMA-Adapter [51] series and LLaVA [22] utilize adapter modules or simple linear layers to merge visual features with LMs, aimed at minimizing the interference between visual and textual data during training while optimizing the model’s capacity to understand and generate visual information.
TABLE II: The five core subtasks of fault diagnosis
@@
II-C Key Tasks of GII
We categorize industrial tasks into four primary types based on their nature and specific requirements: classification, detection/localization, planning & control, and question & answer (Q&A). Each task type presents unique challenges and demands specialized technological approaches, particularly leveraging IIoT, edge computing, and advanced communication technologies. Detailed discussions on the applications and solutions tailored to each task type will be provided in the subsequent sections.
@@ II-C1 Classification
Mechanical Fault Diagnosis. The fault diagnosis tasks map measures signal features relative to the state of mechanical components [13, 12, 13, 52]. This process involves determining the basic condition of whether a fault exists. It may include more complex estimations such as the fault level, the severity of the fault, the remaining useful life, and the specific cause of the fault. Table II details these five essential sub-tasks of fault diagnosis. Equipment fault diagnosis demands high quality and quantity of data, especially when dealing with different types of equipment and their respective fault types, where relevant data are often scarce. Although there are some open-source datasets available, such as CWRU [53], and XJTU-SY [54], which provide some support for model development, they still fail to fully meet the data requirements of practical applications. Fig. 6 illustrates a typical diagnostic data collection device capable of collecting essential operational data from equipment during actual operations.

### v2

# Towards General Industrial Intelligence: A Survey of Continual Large Models in Industrial IoT

@@ Abstract
Industrial AI is transitioning from traditional deep learning models to large-scale transformer-based architectures, with the Industrial Internet of Things (IIoT) playing a pivotal role. IIoT evolves from a simple data pipeline to an intelligent infrastructure, enabling and enhancing these advanced AI systems. This survey explores the integration of IIoT with large models (LMs) and their potential applications in industrial environments. We focus on four primary types of industrial LMs: language-based, vision-based, time-series, and multimodal models. The lifecycle of LMs is segmented into four critical phases: data foundation, model training, model connectivity, and continuous evolution. First, we analyze how IIoT provides abundant and diverse data resources, supporting the training and fine-tuning of LMs. Second, we discuss how IIoT offers an efficient training infrastructure in low-latency and bandwidth-optimized environments. Third, we highlight the deployment advantages of LMs within IIoT, emphasizing IIoT’s role as a connectivity nexus fostering emergent intelligence through modular design, dynamic routing, and model merging to enhance system scalability and adaptability. Finally, we demonstrate how IIoT supports continual learning mechanisms, enabling LMs to adapt to dynamic industrial conditions and ensure long-term effectiveness. This paper underscores IIoT’s critical role in the evolution of industrial intelligence with large models, offering a theoretical framework and actionable insights for future research.
@@
General Industrial Intelligence, Large Language Models, Continual Learning, Industrial Internet of Things, Model Connectivity.
@@ I Introduction
Figure 2: Industrial AI is transitioning from traditional deep learning models to large-scale transformer-based architectures, while the IIoT is also evolving in this process, advancing from a simple data pipeline to an intelligent infrastructure. Our key insight is that IIoT plays a critical role in supporting the full lifecycle of large models, including data foundation, model training, model connectivity, and continuous model evolution.
IIoT serves as a critical framework for industrial digitalization, comprising three core components: sensing, computing, and networking and communication [1, 2, 3, 4]. The sensing layer gathers real-time data through sensors, cameras, and smart devices; the computing layer processes and analyzes data using edge and cloud computing; and the networking and communication layer enables efficient data transmission and inter-device connectivity with low latency and high bandwidth [5, 6, 7]. This architecture not only supports traditional data collection and transmission but also evolves into an intelligent infrastructure, driving significant improvements in production efficiency and system intelligence [8, 9, 10, 11].
As IIoT evolves rapidly, General Industrial Intelligence (GII) [3, 12, 13, 14] emerges as a frontier goal in the industrial domain. The architecture of GII is illustrated in Fig. 1 and comprises four layers: the IIoT infrastructure layer, the foundation layer composed of data and models, the algorithm layer, and the application layer. GII aims to enhance the intelligence and flexibility of industrial systems while emphasizing human-centered design. Its core characteristics include:
  • Continual Learning and Optimization: GII systems continuously learn and optimize in dynamic environments to meet evolving production demands, surpassing traditional task-specific AI systems.
  • Adaptation to Open Environments: GII systems operate effectively in complex and unstructured industrial settings, addressing the rapid changes and personalized demands of future manufacturing processes.
  • Multimodal Interaction: GII systems interact with humans and the physical world through multimodal capabilities, including vision, language, and auditory modalities, significantly improving communication efficiency.
  • Collaborative Intelligence: GII systems facilitate information sharing and knowledge collaboration among machines, production lines, factories, supply chains, and broader industrial ecosystems.
These characteristics underscore the role of GII in driving the deep integration of Industry 5.0 with cyber-physical-social systems [15].
The remarkable advancements in large models (LMs) represent a significant step toward achieving GII. Transformer-based models, including large language models (LLMs) [16, 17, 18], vision models [19, 20], time series models [21, 22], and multimodal models [23, 24], accumulate extensive knowledge through pretraining on large datasets, significantly enhancing the automation and diversity of data processing. For example, GPT-3 demonstrates success in language generation and reasoning tasks [18], while Vision Transformer achieves efficient performance in visual tasks [19]. Models like PandaGPT and LLaMA further expand multimodal interaction capabilities, enabling them to handle more complex tasks in industrial settings [24, 17]. Despite their success, deploying these models in industrial scenarios requires addressing challenges such as dynamic data distributions, task complexity, and user-specific requirements.
TABLE I: Survey comparisons on the lifecycle of large models with IIoT support.
Survey Data Foundation Model Training Model Connectivity Model Evolution
IIoT Testbeds [25] $\checkmark$
AIoT [4] $\checkmark$ $\checkmark$
Federated LLMs [26] $\checkmark$
IoT and AIGC [27] $\checkmark$
AIGC in MEN [28] $\checkmark$
Agents Meet 6G [29] $\checkmark$
Embodied Intelligence [3] $\checkmark$
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
As illustrated in Fig. 2, our key insight is that IIoT transitions from a traditional data pipeline to an intelligent infrastructure, playing a pivotal role in supporting the entire lifecycle of large models. This transformation is reflected in the following aspects:
  • Data Foundation: IIoT serves as the primary source and repository for industrial data. Modern industrial systems generate high-frequency unimodal and multimodal data, including text, images, videos, code, and audio [4]. We further explore how techniques such as data denoising, selection, and generation ensure data quality to support model training and fine-tuning.
  • Model Training: IIoT plays a foundational role in model training by providing low-latency and high-bandwidth network infrastructure. This combination with edge computing effectively supports large-scale distributed training and real-time model fine-tuning, promoting efficient learning and dynamic adaptation.
  • Connectivity: The deployment of large models in IIoT environments benefits from modular design and flexible routing mechanisms. These features significantly enhance system flexibility and adaptability, supporting diverse industrial needs and dynamic application scenarios.
  • Continuous Evolution: IIoT enables large models to maintain long-term effectiveness and adaptability in dynamic industrial environments through continual learning mechanisms. By combining traditional continual learning methods with pre-trained model techniques, IIoT addresses data stream changes and incremental task updates, avoiding catastrophic forgetting and improving model performance.
In summary, IIoT evolves from a traditional data pipeline into an intelligent infrastructure, providing critical support across the entire lifecycle of large models. From data foundation to model training, connectivity, and continuous evolution, IIoT demonstrates its indispensable role. To ensure both relevance and quality, we focus on recent publications from the IEEE Communications Society and major AI conferences over the past five years. This survey systematically analyzes the potential of integrating IIoT with large models, highlighting how this integration drives the transformation of industrial intelligence and offering theoretical frameworks and directions for future research.
I-A Distinguishing Features of This Survey
Our innovations and main contributions are summarized as follows:
$\bullet$ Systematic Analysis from a Full Lifecycle Perspective As shown in Table I, this paper is the first to systematically analyze the key role of IIoT in supporting large models from a full lifecycle perspective. Unlike other surveys that focus on individual stages (e.g., data foundation or model training), this paper comprehensively covers four core aspects: data foundation, model training, model connectivity, and continuous evolution. It reveals how IIoT provides comprehensive support for the development, deployment, and dynamic adaptation of large models.
$\bullet$ Connectivity as a Driver of Emergent Intelligence This paper introduces the innovative concept of “connectivity generates emergent intelligence” and explores the central role of IIoT in model connectivity. Through modular large model design, model routing, and model merging, IIoT enables single devices to support multi-task processing, such as dynamically allocating and combining multiple sub-models to handle complex tasks. This modularity and flexibility not only enhance the intelligence of industrial systems but also significantly expand the application potential of IIoT in dynamic environments, addressing the limitations of existing surveys that overlook the importance of connectivity.
$\bullet$ Expanding the Industrial Applications of Continual Learning This paper deeply examines the critical role of continual learning mechanisms in dynamic industrial environments. It particularly focuses on how the combination of traditional continual learning methods with pre-trained model techniques enhances the long-term adaptability and effectiveness of models. Unlike existing studies that remain theoretical, this paper provides practical solutions tailored to the dynamic data streams and incremental task characteristics of IIoT, advancing the application of continual learning in industrial environments.
@@ II-A Timeline of GII
$\bullet$ Large Models, where models can continuously learn in changing tasks and environments, significantly expanding the boundaries of machine learning applications, marking a step towards higher levels of GII. Representative models include GPT [34] and LLaMA [17].
@@
II-B Large Models
Table II summarizes the key large models for GII, highlighting their features and applications. In Table II, the GPT series models are closed-source, meaning their parameters are neither accessible nor modifiable. Researchers can only guide the model’s predictions by adjusting input data or employing techniques like prompt engineering. In contrast, the remaining models are open-source, allowing access to and modification of model parameters. This enables researchers to fine-tune the models for specific tasks, thereby significantly improving their performance.
TABLE II: Key Large Models for General Industrial Intelligence.
Model Type Model Name Key Features Applications
LLMs GPT, LLaMA Text generation Semantic understanding Industrial fault diagnosis Autonomous driving
Vision Models ViT, CLIP
Visual QA
Image classification
Industrial anomaly detection
Human-robot collaboration
Time Series Models GPT4TS, Time-LLM Time series specific
Equipment predictive maintenance
Energy consumption forecasting
Multimodal Models PandaGPT, BLIP-2 LLaVA, LLaMA-Adapter Multimodal capabilities Industrial scene understanding Industrial multimedia content generation
@@ II-B5 Multimodal Models
Multimodal models are machine learning frameworks designed to process and analyze data from multiple input types, such as text, images, and audio, within a unified architecture. Unlike single-modality models that focus exclusively on one type of data (e.g., text-only or image-only), multimodal models integrate information across multiple modalities. This integration allows these models to leverage complementary features from different data types, resulting in enhanced task accuracy and improved robustness. For example, combining visual and textual information enables a deeper understanding of contextual relationships that would be missed by single-modality approaches. Below are some notable examples:
$\bullet$ BLIP [43] is a multimodal model focused on improving image description and comprehension tasks. By leveraging pre-training techniques like contrastive learning, BLIP effectively aligns textual and visual representations. Its successor, BLIP-2 [44], adopts a dual-stream Transformer architecture that processes visual and textual data streams independently before integrating them with an enhanced cross-modal attention mechanism. This approach significantly improves the alignment between images and text, enabling more accurate image description and understanding.
$\bullet$ MiniGPT-4 [45] is a compact version of GPT tailored for multimodal tasks involving images and text. Despite its reduced scale, it maintains the generative capabilities of the GPT series while focusing on efficient processing of multimodal inputs, making it suitable for real-time applications.
$\bullet$ LLaMA-Adapter [46] and LLaVA [23] employ adapter modules or simple linear layers to integrate visual features into language models. These methods aim to minimize interference between visual and textual data during training while optimizing the model’s ability to understand and generate visual information. This design enables efficient multimodal learning with reduced computational overhead.
TABLE III: The five core subtasks of fault diagnosis.
@@
II-C Key Tasks of Large Model in GII
Building on the introduction to Large Models in Section II-B, this section focuses on the core industrial application capabilities enabled by these LMs. Industrial tasks are categorized into four primary types based on their nature and specific requirements: classification, detection/localization, planning & control, and question & answer (Q&A). Each task type presents unique challenges and demands specialized technological approaches, particularly leveraging IIoT, edge computing, and advanced communication technologies. Detailed discussions on the applications and solutions tailored to each task type will be provided in the subsequent sections.
@@ II-C1 Classification
Mechanical Fault Diagnosis. The fault diagnosis tasks map measures signal features relative to the state of mechanical components [14, 47, 48]. This process involves determining the basic condition of whether a fault exists. It may include more complex estimations such as the fault level, the severity of the fault, the remaining useful life, and the specific cause of the fault. Table III details these five essential sub-tasks of fault diagnosis. Equipment fault diagnosis demands high quality and quantity of data, especially when dealing with different types of equipment and their respective fault types, where relevant data are often scarce. Although there are some open-source datasets available, such as CWRU [49], and XJTU-SY [50], which provide some support for model development, they still fail to fully meet the data requirements of practical applications. Fig. 4 illustrates a typical diagnostic data collection device capable of collecting essential operational data from equipment during actual operations.